### v2

# Quantum Theory and Unusual Dielectric Functions of Graphene

Abstract
We address the spatially nonlocal dielectric functions of graphene at any frequency derived starting from the first principles of thermal quantum field theory using the formalism of the polarization tensor. After a brief review of this formalism, the longitudinal and transverse dielectric functions are considered at any relationship between the frequency and the wave vector. The analytic properties of their real and imaginary parts are investigated at low and high frequencies. Emphasis is given to the double pole at zero frequency which arises in the transverse dielectric function. The role of this unusual property for solving the problem of disagreement between experiment and theory in the Casimir effect is discussed. We guess that a more complete dielectric response of ordinary metals should also be spatially nonlocal and its transverse part may possess the double pole in the region of evanescent waves.
I Introduction
It is common knowledge that quantum physics originated in 1900 from the work of Max Planck who derived the distribution law for the monochromatic radiation, introduced a concept of the quantum of energy and the new fundamental constant $h$ now known as the Planck constant 1I . In 1905, Albert Einstein arrived at a concept of the quanta of light, which were later called photons, and explained on this basis the photoelectric effect 2I . However, the first quantum theory, quantum mechanics, was created one hundred years ago by Werner Heisenberg (1925) 3I and Erwin Schödinger (1926) 4I . The relativistic quantum mechanics formulated by Paul Dirac in 1928 5I introduced the concept of antiparticles and opened a way to the formulation of quantum field theory and its applications to all fundamental interactions of nature during the twentieth century.
For a long time, it was believed that quantum theory is only needed to describe very small submicroscopic objects on atomic and even subatomic scales. Later it was understood, however, that there are a lot of macroscopic quantum phenomena characterized by the scales greatly exceeding the atomic ones. These are the superconductivity, superfluidity, quantum Hall and Josephson effects, Bose-Einstein condensation, the Casimir effect etc. Currently the macroscopic quantum phenomena are not of only an academic interest, but are widely used in many technological and industrial applications, as well as in metrology.
Considerable recent attention has been focussed on various novel materials described by quantum theory, and the two-dimensional sheet of carbon atoms called graphene occupies a prominent place among them. Graphene is remarkable for many reasons including its unique mechanical, electrical and optical properties 1 ; 2 ; 3 . At energies below approximately 3 eV, graphene is well described by the Dirac model. This means that the field of either massless or very light electronic quasiparticles in graphene satisfies the (2+1)-dimensional Dirac equation rather than the Schrödinger equation which describes the standard quasiparticles considered in condensed matter physics. In doing so, the Fermi velocity in graphene $v_{F}\approx c/300$ plays the same role as the speed of light $c$ in the usual Dirac equation.
The response of graphene to the electromagnetic field is spatially nonlocal. It is commonly described by the tensors of electric conductivity or dielectric permittivity. For a graphene sheet in the absence of constant magnetic field, these tensors are characterized by the two functions each depending on frequency $\omega$, the two-dimensional wave vector $\bf q$} and on temperature $T$ (for the gapped and doped graphene they also depend on the mass gap parameter and chemical potential). These are the longitudinal $\sigma^{\rm L}$ and transverse $\sigma^{\rm T}$ conductivities and the corresponding dielectric functions $\varepsilon^{\rm L}$ and $\varepsilon^{\rm T}$. In the case of two spatial dimensions, the dielectric functions are expressed in terms of conductivities as 4 ; 5
$\varepsilon^{\rm{L,T}}(\omega,\mbox{$q$})=1+\frac{2\pi iq}{\omega}\sigma^{\rm{L,T}}(\omega,\mbox{$q$}),$ (1)
where $q=|\mbox{$q$}|=\sqrt{q_{1}^{2}+q_{2}^{2}}$. This means that in the Gaussian units used here and below the conductivities of graphene have the dimension of cm/s (for the three-dimensional materials, the dimension of conductivity is 1/s).
The response functions of graphene were investigated in both the spatially local ($\mbox{$q$}=0,~\varepsilon^{\rm L}=\varepsilon^{\rm T}=\varepsilon$) and nonlocal cases using a number of more or less phenomenological approaches such as the hydrodynamic model 5a ; 5b ; 5c , Boltzmann transport theory and the Drude model 11 ; 11a ; 11b ; 13 ; 14 ; 18 ; 24 ; 25 ; 31 , current-current correlation functions and the random phase approximation 9 ; 12 ; 12a ; 18a ; 21 ; 23a ; 26 ; 26a ; 26b ; 29a , density-functional theory 23 ; 23aa , Kubo response theory 6 ; 7 ; 7a ; 7ab ; 8 ; 15 ; 16 ; 17 ; 27 ; 27a ; 28 ; 29 , modeling graphene optics in terms of Lorentz-type oscillators 31a and by using the Fresnel reflection coefficients 32 , etc. (see also the reviews 34 ; 35 ; 36 ). The obtained results are of different levels of accuracy and areas of application. In the framework of the Dirac model, however, the response functions of graphene can be found exactly starting from the first principles of thermal quantum field theory using the formalism of the polarization tensor in (2+1)-dimensions.
In this article, we discuss the properties of the longitudinal and transverse dielectric functions of graphene expressed in the framework of the quantum field theory via the components of the polarization tensor. The dependence of these functions on frequency is investigated over the entire region of positive frequencies (the dependence of the dielectric function of graphene on temperature is investigated in Ref. 39a ). It is shown that these functions possess some usual properties characteristic of common materials. Thus, they satisfy the Kramers-Kronig relations, go to unity with an indefinitely increasing frequency, and have the positive imaginary parts as it must be in accordance with the second law of thermodynamics 37 . At the same time, we demonstrate that the transverse dielectric function of graphene possesses an unusual property by having the double pole at zero frequency (it is generally believed that at zero frequency the response functions of metallic and dielectric materials have a single pole and are regular, respectively). We propose that in the region of evanescent waves the transverse dielectric function of ordinary metals may have the double pole as well.
We start in Section 2 with a brief review of the most necessary results regarding the polarization tensor of graphene. Then in Section 3 we consider the properties of dielectric functions of graphene expressed via the polarization tensor at low frequencies. Section 4 is devoted to the case of high frequencies. Sections 5 and 6 contain the discussion and our conclusions. For the sake of clarity in presentation, all mathematical equations are written for the case of a pristine graphene possessing the zero mass gap parameter and chemical potential. However, all the results presented below remain valid for the gapped and doped graphene sheets.
II Polarization tensor of graphene
In the one-loop approximation, the interaction of electronic quasiparticles in graphene with the electromagnetic field is described by the quasiparticle loop diagram having two photon legs. It is represented by the polarization tensor $\Pi_{\mu\nu}$, where $\mu,\nu=0,1,2$. At zero temperature, the polarization tensor has long been calculated within (2+1)-dimensional quantum field theory 38 ; 39 . Specifically for graphene, whose properties are temperature-dependent, the polarization tensor was studied in detail at both zero and nonzero temperature 6 ; 40 ; 41 ; 42 ; 43 ; 44 . In the latter case, the formalism of thermal quantum field theory in the Matsubara formulation has been used.
The expressions for the polarization tensor of graphene valid over the entire plane of complex frequency, including the real frequency axis, were obtained in 45 ; 46 (the previously obtained expressions 44 are valid only at the pure imaginary Matsubara frequencies). They were used for investigation of the electric conductivity 47 ; 48 ; 49 ; 50 and reflectivity 51 ; 52 ; 53 of graphene, as well as of the Casimir and Casimir-Polder forces in out-of-thermal-equilibrium graphene systems 54 ; 55 ; 56 ; 57 ; 58 ; 59 .
All components of the polarization tensor can be expressed via the two independent quantities 44 , e.g., via $\Pi_{00}$ and
$\Pi(\omega,\mbox{$q$})=q^{2}\Pi_{\mu}^{\,\mu}(\omega,\mbox{$q$})+\left(\frac{\omega^{2}}{c^{2}}-q^{2}\right)\Pi_{00}(\omega,\mbox{$q$}).$ (2)
There are different but mathematically equivalent representations for the quantities $\Pi_{00}$ and $\Pi$. Below we use that ones presented in Ref.  54 .
Note that the entire range of positive frequencies from zero to infinity can be divided into the regions of evanescent, $0<\omega<cq$, and propagating, $\omega\geqslant cq$, waves. In its turn, in the region of evanescent waves, it is convenient to separate the subregion of strongly evanescent waves, $0<\omega<v_{F}q$. The explicit expressions for $\Pi_{00}$ and $\Pi$ have different forms in the regions $0<\omega<v_{F}q$ and $\omega>v_{F}q$. We start with the region $0<\omega<v_{F}q$, i.e., with the strongly evanescent waves. In this region, for $\Pi_{00}$ one has 54
[…]
$\left.+\frac{1}{2\sqrt{v_{F}^{2}q^{2}-\omega^{2}}}\left[\int\limits_{0}^{v_{F}q-\omega}\!\!\!\!\frac{dx}{e^{\,\beta x}+1}F_{1}(x)-\!\!\!\int\limits_{0}^{v_{F}q+\omega}\!\!\!\!\frac{dx}{e^{\,\beta x}+1}F_{2}(x)\right]\right\}$
and
${\rm Im}\,\Pi_{00}(\omega,\mbox{$q$})=\frac{4\alpha\hbar c}{v_{F}^{2}\sqrt{v_{F}^{2}q^{2}-\omega^{2}}}\left[\,\,\int\limits_{v_{F}q-\omega}^{\infty}\!\!\!\!\frac{dx}{e^{\,\beta x}+1}F_{3}(x)-\!\!\int\limits_{v_{F}q+\omega}^{\infty}\!\!\!\!\frac{dx}{e^{\,\beta x}+1}F_{4}(x)\right],$ (4)
where $\alpha=e^{2}/(\hbar c)$ is the fine structure constant, $\beta=\hbar/(2k_{B}T)$, $k_{B}$ being the Boltzmann constant, and
$F_{1,2}(x)=\sqrt{v_{F}^{2}q^{2}-(x\pm\omega)^{2}},\quad F_{3,4}(x)=\sqrt{(x\pm\omega)^{2}-v_{F}^{2}q^{2}}.$
In a similar way, for $\Pi$ one obtains 54
[…]
$\hskip-14.22636pt\left.+\frac{\sqrt{v_{F}^{2}q^{2}-\omega^{2}}}{2}\left[\int\limits_{0}^{v_{F}q-\omega}\!\!\!\!\!\frac{dx}{e^{\,\beta x}+1}\frac{(x+\omega)^{2}}{F_{1}(x)}-\!\!\!\!\int\limits_{0}^{v_{F}q+\omega}\!\!\!\!\!\frac{dx}{e^{\,\beta x}+1}\frac{(x-\omega)^{2}}{F_{2}(x)}\right]\right\}$
and
${\rm Im}\,\Pi(\omega,\mbox{$q$})=\frac{4\alpha\hbar}{v_{F}^{2}c}\sqrt{v_{F}^{2}q^{2}-\omega^{2}}\left[\,\,\int\limits_{v_{F}q+\omega}^{\infty}\!\!\!\!\frac{dx}{e^{\,\beta x}+1}\frac{(x-\omega)^{2}}{F_{4}(x)}-\!\!\!\int\limits_{v_{F}q-\omega}^{\infty}\!\!\!\!\frac{dx}{e^{\,\beta x}+1}\frac{(x+\omega)^{2}}{F_{3}(x)}\right].$ (6)
In the remaining region of evanescent waves $v_{F}q<\omega<cq$ and in the region of propagating waves $\omega\geqslant cq$, the quantities $\Pi_{00}$ and $\Pi$ are given by the unified expressions 54 . Thus, for $\Pi_{00}$ one has
[…]
$\hskip-14.22636pt\left.\times\left[\int\limits_{0}^{\infty}\!\!\frac{dx}{e^{\,\beta x}+1}F_{3}(x)-\!\!\!\!\int\limits_{\omega+v_{F}q}^{\infty}\!\!\!\!\frac{dx}{e^{\,\beta x}+1}F_{4}(x)+\!\!\!\!\int\limits_{0}^{v_{F}q-\omega}\!\!\!\!\frac{dx}{e^{\,\beta x}+1}F_{4}(x)\right]\right\}$
and
${\rm Im}\,\Pi_{00}(\omega,\mbox{$q$})=\frac{\alpha\hbar c}{\sqrt{\omega^{2}-v_{F}^{2}q^{2}}}\left[\pi q^{2}-\frac{4}{v_{F}^{2}}\!\!\int\limits_{\,-v_{F}q}^{v_{F}q}\!\!\!\!dx\frac{\sqrt{v_{F}^{2}q^{2}-x^{2}}}{e^{\beta(\omega+x)}+1}\right].$ (8)
For $\Pi$ the following expressions are valid
[…]
$\times\left[\int\limits_{0}^{\infty}\!\!\frac{dx}{e^{\,\beta x}+1}\frac{(x+\omega)^{2}}{F_{3}(x)}-\!\!\int\limits_{\omega+v_{F}q}^{\infty}\!\!\!\!\frac{dx}{e^{\,\beta x}+1}\frac{(x-\omega)^{2}}{F_{4}(x)}\right.\left.\left.+\!\!\int\limits_{0}^{\omega-v_{F}q}\!\!\!\!\frac{dx}{e^{\,\beta x}+1}\frac{(x-\omega)^{2}}{F_{4}(x)}\right]\right\}$ (9)
and
${\rm Im}\,\Pi(\omega,\mbox{$q$})=\frac{\alpha\hbar}{v_{F}^{2}c}\sqrt{\omega^{2}-v_{F}^{2}q^{2}}\left[-\pi v_{F}^{2}q^{2}+4\!\!\int\limits_{-v_{F}q}^{v_{F}q}\!\!\!\!\frac{dx}{e^{\beta(\omega+x)}+1}\frac{x^{2}}{\sqrt{v_{F}^{2}q^{2}-x^{2}}}\right].$ (10)
The polarization tensor is gauge-invariant and, as a consequence, satisfies the transversality condition 38 ; 39 ; 40 ; 41 ; 42 ; 43 ; 44 ; 45 ; 46
$q^{\mu}\Pi_{\mu\nu}(\omega,\mbox{$q$})=0,\,\,\,q^{\mu}=(\omega/c,q^{1},q^{2}).$ (11)
Now we use an expression for the current arising due an application of the electromagnetic field
$J^{\mu}(\omega,\mbox{$q$})=\frac{c}{4\pi\hbar}\Pi^{\mu\nu}(\omega,\mbox{$q$})A_{\nu}(\omega,\mbox{$q$}),$ (12)
where $A_{\nu}$ is the vector potential, and the microscopic relativistically covariant Ohm’s law 62a
$J^{\mu}(\omega,\mbox{$q$})=\sigma^{\mu\nu}(\omega,\mbox{$q$})E_{\nu}(\omega,\mbox{$q$}),$ (13)
where $E_{\nu}$ is the 3-vector of the electric field (see also Ref. 62b for a definition of the relativistically covariant vectors of electric and magnetic fields). By employing Eqs. (12), (13) and $E_{\nu}=(i\omega/c)A_{\nu}$, one expresses the tensor of electric conductivity via the polarization tensor 9 ; 60 ; 61 ; 62 ; 63 ; 64
$\sigma^{\mu\nu}(\omega,\mbox{$q$})=\frac{c^{2}}{4\pi\hbar}\frac{\Pi^{\mu\nu}(\omega,\mbox{$q$})}{i\omega}.$ (14)
With the help of this equation, the longitudinal and transverse conductivities of graphene are presented in the form 47 ; 48 ; 49 ; 50
$\sigma^{\rm L}(\omega,\mbox{$q$})=-\frac{i\omega}{4\pi\hbar q^{2}}\Pi_{00}(\omega,\mbox{$q$}),\qquad\sigma^{\rm T}(\omega,\mbox{$q$})=\frac{ic^{2}}{4\pi\hbar q^{2}\omega}\Pi(\omega,\mbox{$q$}).$ (15)
Finally, using Eq. (1), for the longitudinal and transverse dielectric functions of graphene we obtain 47 ; 65
$\varepsilon^{\rm L}(\omega,\mbox{$q$})=1+\frac{1}{2\hbar q}\Pi_{00}(\omega,\mbox{$q$}),\qquad\varepsilon^{\rm T}(\omega,\mbox{$q$})=1-\frac{c^{2}}{2\hbar q\omega^{2}}\Pi(\omega,\mbox{$q$}).$ (16)
Note that recently the basics of quantum field theoretical approach to a description of the electric conductivity and dielectric response of graphene were cast under doubt. It was noticed 66 that some of the results obtained using quantum field theory are in disagreement with those following from the Kubo model. Based on the Kubo formula, the polarization tensor $\Pi^{\mu\nu}$ in Refs.  (12) and (14) was replaced with the so-called ”regularized” quantity $\widetilde{\Pi}^{\mu\nu}$ defined as 66
$\widetilde{\Pi}^{\mu\nu}(\omega,\mbox{$q$})={\Pi}^{\mu\nu}(\omega,\mbox{$q$})-\lim\limits_{\omega\to 0}{\Pi}^{\mu\nu}(\omega,\mbox{$q$}).$ (17)
It was shown, however, that the polarization tensor $\Pi^{\mu\nu}$ is defined uniquely and cannot be modified with no violation of first principles of quantum theory 67 . The derivation of Eq. (15) with $\widetilde{\Pi}_{\mu\nu}(\omega,\mbox{$q$})$ in place of ${\Pi}_{\mu\nu}(\omega,\mbox{$q$})$ from the Kubo formula in Ref. 66 used the nonrelativistic concept of causality rather than the relativistic one as would be correct for the Dirac model. Specifically, Ref. 66 applied the one-sided Fourier transforms from 0 to $\infty$ instead of the two-sided one from -$\infty$ to $\infty$, which must be used in the relativistic theory, and obtained the subtracted term in Eq. (17) making an integration by parts in the integral from 0 to $\infty$. This resulted in a violation of the gauge invariance and in other physically unacceptable consequences 68 .
Thus, there is no any contradiction between the results obtained using the quantum field theory and the Kubo model if the latter is applied correctly. When using the two-sided Fourier transforms, as one should do in application to the relativistic systems, such as graphene, the Kubo formula results in the correct Eq. (14) with the polarization tensor $\Pi^{\mu\nu}$ 68 .
III Dielectric functions of graphene at low frequencies
Here, we consider the properties of both the longitudinal and transverse dielectric functions of graphene in the region of strongly evanescent waves $0<\omega<v_{F}q$. Substituting Eqs. (3) and (4) in the first equality of Eq. (16), the real and imaginary parts of the longitudinal dielectric function are found in the form
[…]
$\left.+\frac{1}{2\sqrt{v_{F}^{2}q^{2}-\omega^{2}}}\left[\int\limits_{0}^{v_{F}q-\omega}\!\!\!\!\frac{dx}{e^{\,\beta x}+1}F_{1}(x)-\!\!\!\int\limits_{0}^{v_{F}q+\omega}\!\!\!\!\frac{dx}{e^{\,\beta x}+1}F_{2}(x)\right]\right\}$
and
${\rm Im}\,\varepsilon^{\rm L}(\omega,\mbox{$q$})=\frac{2\alpha c}{v_{F}^{2}q\sqrt{v_{F}^{2}q^{2}-\omega^{2}}}\left[\,\,\int\limits_{v_{F}q-\omega}^{\infty}\!\!\!\!\frac{dx}{e^{\,\beta x}+1}F_{3}(x)-\!\!\int\limits_{v_{F}q+\omega}^{\infty}\!\!\!\!\frac{dx}{e^{\,\beta x}+1}F_{4}(x)\right].$ (19)
As is seen in Eqs. (18) and (19), in the limiting case $\omega\to 0$ one has
$\lim\limits_{\omega\to 0}{\rm Re}\,\varepsilon^{\rm L}(\omega,\mbox{$q$})=1+\frac{\pi\alpha c}{2v_{F}}+\frac{8\alpha c\ln 2}{v_{F}^{2}q}\,\frac{k_{B}T}{\hbar},$
$\lim\limits_{\omega\to 0}{\rm Im}\,\varepsilon^{\rm L}(\omega,\mbox{$q$})=0,$ (20)
i.e., the longitudinal dielectric function of graphene is regular at zero frequency. From Eq.  (19) it is seen that ${\rm Im}\,\varepsilon^{\rm L}(\omega,\mbox{$q$})>0$ because the integrand in the first integral is larger than in the second and integrated over the wider interval.
As an example, in Figure 1, ${\rm Re}\,\varepsilon^{\rm L}$ given by Eq. (18) is shown at T=300T=300K,
[…]
Figure 1: The magnitude of real part of the longitudinal dielectric function of graphene is shown versus frequency for $T=300$ K and $q=100~\mbox{cm}^{-1}$ in the logarithmic scale. The threshold at $\omega=v_{F}q$ is marked by the dashed vertical line.
Figure. 2 shows the imaginary part of the longitudinal response function of graphene, ${\rm Im}\,\varepsilon^{\rm L}$, given by Eq.  (19) at T=300T=300K, $q=100~{\rm cm}^{-1}$ as the function of frequency in the region to the left of the vertical dashed line $\omega=v_{F}q$. As is seen in Figure 2, ${\rm Im}\,\varepsilon^{\rm L}$ goes to infinity when $\omega$ approaches $v_{F}q$ from the left, as it does ${\rm Re}\,\varepsilon^{\rm L}$ .
Figure 2: The imaginary part of the longitudinal dielectric function of graphene is shown versus frequency for $T=300$ K and $q=100~\mbox{cm}^{-1}$ in the logarithmic scale. The threshold at $\omega=v_{F}q$ is marked by the dashed vertical line.
The real and imaginary parts of the transverse dielectric function of graphene at low frequencies are obtained by substituting Eqs. (5) and (6) to the second equation in (16). The result is
$\hskip-17.07164pt{\rm Re}\,\varepsilon^{\rm T}(\omega,\mbox{$q$})=1-\frac{\pi\alpha cq}{2\omega^{2}}\sqrt{v_{F}^{2}q^{2}-\omega^{2}}-\frac{4\alpha c}{v_{F}^{2}q}\,\frac{\ln 2}{\beta}$ (21)
$\hskip-17.07164pt-\frac{2\alpha c\sqrt{v_{F}^{2}q^{2}-\omega^{2}}}{v_{F}^{2}q\omega^{2}}\left[\int\limits_{0}^{v_{F}q-\omega}\!\!\!\!\!\frac{dx}{e^{\,\beta x}+1}\frac{(x+\omega)^{2}}{F_{1}(x)}-\!\!\!\!\int\limits_{0}^{v_{F}q+\omega}\!\!\!\!\!\frac{dx}{e^{\,\beta x}+1}\frac{(x-\omega)^{2}}{F_{2}(x)}\right]$
and
${\rm Im}\,\varepsilon^{\rm T}(\omega,\mbox{$q$})=\frac{2\alpha c}{v_{F}^{2}q\omega^{2}}\sqrt{v_{F}^{2}q^{2}-\omega^{2}}\left[\,\,\int\limits_{v_{F}q-\omega}^{\infty}\!\!\!\!\frac{dx}{e^{\,\beta x}+1}\frac{(x+\omega)^{2}}{F_{3}(x)}-\!\!\!\int\limits_{v_{F}q+\omega}^{\infty}\!\!\!\!\frac{dx}{e^{\,\beta x}+1}\frac{(x-\omega)^{2}}{F_{4}(x)}\right].$ (22)
[…]
The behavior of the last term in Eq. (21) in the limiting case $\omega\to 0$ is not so evident. To determine it, we introduce the parameter $\kappa\equiv\omega/(v_{F}q)$ and perform the changes of variables $t=x/(v_{F}q)\pm\kappa$ in the first and second integrals in the squared brackets, respectively. Then one obtains
$\int\limits_{0}^{v_{F}q-\omega}\!\!\!\!\!\frac{dx}{e^{\,\beta x}+1}\frac{(x+\omega)^{2}}{F_{1}(x)}-\!\!\!\!\int\limits_{0}^{v_{F}q+\omega}\!\!\!\!\!\frac{dx}{e^{\,\beta x}+1}\frac{(x-\omega)^{2}}{F_{2}(x)}=v_{F}^{2}q^{2}(I_{1}+I_{2}),$ (23)
where
$I_{1}=\int\limits_{\kappa}^{1}\!\!\frac{t^{2}dt}{\sqrt{1-t^{2}}}\left(\frac{1}{e^{\gamma t}e^{-\gamma\kappa}+1}-\frac{1}{e^{\gamma t}e^{\gamma\kappa}+1}\right),$
$I_{2}=\int\limits_{-\kappa}^{\kappa}\frac{dt}{e^{\gamma t}e^{\gamma\kappa}+1}\,\frac{t^{2}}{\sqrt{1-t^{2}}},\quad\gamma\equiv v_{F}q\beta.$ (24)
Under the condition $\omega\ll v_{F}q$, i.e. $\kappa\ll 1$, at fixed $q,\,T\neq 0$, we expand the integrand in $I_{1}$ in powers of the small parameter $\beta\kappa$ and obtain
$I_{1}=\gamma\kappa B_{1}+O(\gamma^{3}\kappa^{3}),\quad B_{1}\equiv 2\int\limits_{0}^{1}\!\!\!\frac{t^{2}dt}{\sqrt{1-t^{2}}}\frac{e^{\gamma t}}{(e^{\gamma t}+1)^{2}}.$ (25)
By making similar expansion in $I_{2}$, one finds that in the lowest order $I_{2}=\kappa^{3}/3$ and, thus, it does not contribute to the behavior of ${\rm Re}\,\varepsilon^{\rm T}$ at low frequencies.
As a result, substituting Eqs. (23) and (25) to Eq. (21), for the low-frequency behavior of the real part of transverse dielectric function of graphene one finds
${\rm Re}\,\varepsilon^{\rm T}(\omega,\mbox{$q$})=1-8\ln 2\frac{\alpha c}{v_{F}^{2}q}\,\frac{k_{B}T}{\hbar}-\frac{\alpha\hbar cv_{F}q^{2}}{k_{B}T}\,\frac{B_{1}}{\omega}-\frac{\pi\alpha cv_{F}q^{2}}{2\omega^{2}}.$ (26)
Note that the ”regularized” polarization tensor (17) was introduced in Ref. 66 with the aim to remove the last term in Eq. (26) which was considered by the authors of Ref. 66 as ”nonphysical”. The presence of this term, however, is in agreement with all physical principles and was confirmed experimentally by measuring the Casimir force in graphene systems 69 ; 70 . Using the current-current correlation functions 21 and the polarization tensor 44 , it was predicted that in the systems with a graphene layer at nonzero temperature the Casimir force reaches the high-temperature asymptotics equal to one-half of that valid for ideal metals already at short separations. According to Ref. 21 , for graphene the high-temperature regime is reached under the condition $ak_{B}T\gtrsim\hbar v_{F}\approx 0.0033\hbar c$. Using the formalism of the polarization tensor, Ref. 44 indicates the application condition of the high-temperature regime for the system of a pristine graphene sheet parallel to a metallic plate as $ak_{B}T\gg\alpha\ln\alpha^{-1}\hbar c/[2\zeta(3)]\approx 0.015\hbar c$, where $\zeta(z)$ is the Riemann zeta function (note that Ref. 44 uses the units with $\hbar=c=k_{B}=1$). Employing the more exact asymptotic expressions for the polarization tensor, it was shown 68j that for this system the high-temperature regime takes place under a less severe condition
$ak_{B}T\gg\frac{\hbar v_{F}^{2}}{8\ln 2~\alpha c}\approx 0.000276\hbar c.$ (27)
This condition is in numerical agreement with that of Ref. 21 , but, according to the condition of Ref. 44 , the high-temperature regime begins at much larger values of $ak_{B}T$. The results of numerical computations 68j are in agreement with the application condition 21 of the high-temperature regime. The resulting unusually big finite-temperature Casimir effect in graphene systems calculated using the polarization tensor was measured in Refs. 69 ; 70 .
Now we consider the imaginary part of $\varepsilon^{\rm T}$ given by Eq. (22). By performing the same changes of variables as above, $t=x/(v_{F}q)\pm\kappa$, in the first and second integrals in the square brackets, respectively, and expanding in powers of the small parameter $\beta\kappa$ like this was done in the integral $I_{1}$ in Eq. (24), we obtain
$\int\limits_{v_{F}q-\omega}^{\infty}\!\!\!\!\frac{dx}{e^{\,\beta x}+1}\frac{(x+\omega)^{2}}{F_{3}(x)}-\!\!\!\int\limits_{v_{F}q+\omega}^{\infty}\!\!\!\!\frac{dx}{e^{\,\beta x}+1}\frac{(x-\omega)^{2}}{F_{4}(x)}$
$=v_{F}^{2}q^{2}\int\limits_{1}^{\infty}\!\!\frac{t^{2}dt}{\sqrt{1-t^{2}}}\left(\frac{1}{e^{\gamma t}e^{-\gamma\kappa}+1}-\frac{1}{e^{\gamma t}e^{\gamma\kappa}+1}\right)=v_{F}^{2}q^{2}\left[\gamma\kappa B_{2}+O(\gamma^{3}\kappa^{3})\right],$ (28)
where
$B_{2}\equiv 2\int\limits_{1}^{\infty}\!\!\!\frac{t^{2}dt}{\sqrt{t^{2}-1}}\frac{e^{\gamma t}}{(e^{\gamma t}+1)^{2}}.$ (29)
Substituting Eq. (28) in Eq. (22), the low-frequency behavior of the imaginary part of transverse dielectric function of graphene takes the form
${\rm Im}\,\varepsilon^{\rm T}(\omega,\mbox{$q$})=\frac{\alpha\hbar cv_{F}q^{2}}{k_{B}T}\,\frac{B_{2}}{\omega}.$ (30)
It is seen that the second line of Eq. (28) is evidently positive and, thus, ${\rm Im}\,\varepsilon^{\rm T}>0$ as it should be.
In Figure 3, $|{\rm Re}\,\varepsilon^{\rm T}|$ given by Eq. (21) is shown as the function of frequency in the region $\omega<v_{F}q$ for the same values of $T$ and $q$ as in Figures 1 and 2. When $\omega$ approaches $v_{F}q$, ${\rm Re}\,\varepsilon^{\rm T}$ approaches the negative constant. The asymptotic expression (26) is well applicable at all $\omega<v_{F}q$.
Figure 3: The magnitude of real part of the transverse dielectric function of graphene is shown versus frequency for $T=300$ K and $q=100~\mbox{cm}^{-1}$ in the logarithmic scale. The threshold at $\omega=v_{F}q$ is marked by the dashed vertical line.
Figure 4 shows the imaginary part of the transverse response function of graphene, ${\rm Im}\,\varepsilon^{\rm T}$, given by Eq. (22) at T=300T=300~K, $q=100~\mbox{cm}^{-1}$ as the function of frequency in the region to the left of the vertical line $\omega=v_{F}q$ . As is seen in Figure 4, ${\rm Im}\,\varepsilon^{\rm T}$ goes to zero when $\omega$ approaches $v_{F}q$ from the left. The asymptotic expression (30) is well applicable for $\omega<2\times 10^{9}$ rad/s.
Figure 4: The imaginary part of the transverse dielectric function of graphene is shown versus frequency for $T=300$ K and $q=100~\mbox{cm}^{-1}$ in the logarithmic scale. The threshold at $\omega=v_{F}q$ is marked by the dashed vertical line.
As opposed to Eq. (20) for $\varepsilon^{\rm L}$, where one can consider the limit of zero $T$, Eqs. (26) and (30) for $\varepsilon^{\rm T}$ are obtained under a condition $T\neq 0$. The exact expressions for $\varepsilon^{\rm T}$ for any $\omega$ at $T=0$ are 71
${\rm Im}\,\varepsilon^{\rm T}(\omega,\mbox{$q$})=\left\{\begin{array}[]{cl}-\frac{\pi\alpha qc}{\omega^{2}}\,\sqrt{v_{F}^{2}q^{2}-\omega^{2}},&\omega<v_{F}q,\\[2.84526pt] i\frac{\pi\alpha qc}{2\omega^{2}}\,\sqrt{\omega^{2}-v_{F}^{2}q^{2}},&\omega>v_{F}q.\end{array}\right$. (31)
Thus, at zero temperature the double pole at zero frequency in $\varepsilon^{\rm T}$ is preserved. Note that for graphene the limiting transitions of $\omega$ and $T$ to zero are not interchangeable.
IV Dielectric functions of graphene at high frequencies
Now we consider the longitudinal and transverse dielectric functions of graphene at all frequencies satisfying the condition $\omega>v_{F}q$. This includes the region of evanescent waves $v_{F}q<\omega<cq$ and the region of propagating waves $\omega\geqslant cq$.
The real and imaginary parts of the longitudinal dielectric function are obtained by substituting Eqs. (7) and (8) in the first equality of Eq. (16)
[…]
$\hskip-14.22636pt\left.\times\left[\int\limits_{0}^{\infty}\!\!\frac{dx}{e^{\,\beta x}+1}F_{3}(x)-\!\!\!\!\int\limits_{\omega+v_{F}q}^{\infty}\!\!\!\!\frac{dx}{e^{\,\beta x}+1}F_{4}(x)+\!\!\!\!\int\limits_{0}^{v_{F}q-\omega}\!\!\!\!\frac{dx}{e^{\,\beta x}+1}F_{4}(x)\right]\right\}$
and
${\rm Im}\,\varepsilon^{\rm L}(\omega,\mbox{$q$})=\frac{\alpha c}{2q\sqrt{\omega^{2}-v_{F}^{2}q^{2}}}\left[\pi q^{2}-\frac{4}{v_{F}^{2}}\!\!\int\limits_{\,-v_{F}q}^{v_{F}q}\!\!\!\!dx\frac{\sqrt{v_{F}^{2}q^{2}-x^{2}}}{e^{\beta(\omega+x)}+1}\right].$ (33)
We consider first the limiting value of ${\rm Re}\,\varepsilon^{\rm L}$ when $\omega\to\infty$. By introducing the new integration variable $y=x-\omega$ in the second integral of Eq. (32), we obtain
$\int\limits_{\omega+v_{F}q}^{\infty}\!\!\!\!\frac{dx}{e^{\,\beta x}+1}F_{4}(x)=-\!\!\!\int\limits_{v_{F}q}^{\infty}\!\!\ \frac{dy}{e^{\beta(\omega+y)}+1}\sqrt{y^{2}-v_{F}^{2}q^{2}},$ (34)
which goes to zero exponentially fast when $\omega\to\infty$ and can be omitted. In the remaining terms in the figure brackets of Eq. (32), we change the integration variable according to $x=\omega y$, introduce the small parameter $\delta=v_{F}q/\omega$ and in the limit $\omega\to\infty$ obtain
$\hskip-19.91692pt\frac{2\ln 2}{\beta}-\frac{1}{\sqrt{\omega^{2}-v_{F}^{2}q^{2}}}\int\limits_{0}^{\infty}\!\!\frac{dx}{e^{\,\beta x}+1}[F_{3}(x)+F_{4}(x)]$ (35)
$\hskip-19.91692pt=\frac{2\ln 2}{\beta}-\frac{\omega}{\sqrt{1-\delta^{2}}}\int\limits_{0}^{\infty}\!\!dy\frac{\sqrt{(y+1)^{2}-\delta^{2}}+\sqrt{(y-1)^{2}-\delta^{2}}}{e^{\,\beta\omega y}+1}=\frac{2\ln 2}{\beta}-2\omega\int\limits_{0}^{\infty}\!\!\frac{dy}{e^{\,\beta\omega y}+1}=0.$
Here, we took into account that when $\omega\to\infty$ it holds $\beta\omega\gg 1$ as well. As a result, the dominant contribution to the integrals is given by $y\ll 1$, so that $(y+1)^{2}\approx(y-1)^{2}\approx 1$ and one can expand the square roots in powers of the small parameter $\delta^{2}$. Thus, in the limiting case $\omega\to\infty$, one obtains
$\lim\limits_{\omega\to\infty}{\rm Re}\,\varepsilon^{\rm L}(\omega,\mbox{$q$})=1,\quad\lim\limits_{\omega\to\infty}{\rm Im}\,\varepsilon^{\rm L}(\omega,\mbox{$q$})=0.$ (36)
The latter equality is an evident consequence of Eq. (33).
From Eq. (33) it follows also that ${\rm Im}\,\varepsilon^{\rm L}>0$. The point is that the integrand in this equation is the decreasing function of $\beta$. Thus, it takes the maximum value for $\beta=0$ (i.e., for $T=\infty$). Then one has
$\max\!\!\int\limits_{-v_{F}q}^{v_{F}q}\!\!\!\!dx\frac{\sqrt{v_{F}^{2}q^{2}-x^{2}}}{e^{\beta(\omega+x)}+1}=\frac{1}{2}\!\int\limits_{-v_{F}q}^{v_{F}q}\!\!\!\!dx\sqrt{v_{F}^{2}q^{2}-x^{2}}=\frac{\pi}{4}v_{F}^{2}q^{2}.$ (37)
Substituting this in Eq. (33), we find that $\min{\rm Im}\,\varepsilon^{\rm L}=0$ and conclude that at all $T<\infty$ it holds ${\rm Im}\,\varepsilon^{\rm L}>0$.
In the domains of Figures 1 and 2 to the right of the dashed vertical lines ($\omega>v_{F}q$), $|{\rm Re}\,\varepsilon^{\rm L}|$ and ${\rm Im}\,\varepsilon^{\rm L}$, respectively, given by Eqs. (32) and (33) are shown as the functions of frequency. When $\omega$ increases from $v_{F}q$ to $1.54\times 10^{12}$, ${\rm Re}\,\varepsilon^{\rm L}$ varies from minus infinity to zero. With further increase of $\omega$, ${\rm Re}\,\varepsilon^{\rm L}$ changes its sign and increases to unity. As to ${\rm Im}\,\varepsilon^{\rm L}$, it abruptly decays to zero at $\omega>v_{F}q$.
The real and imaginary parts of the transverse dielectric function of graphene under the condition $\omega>v_{F}q$ are found from Eqs. (9), (10) and the second equality of Eq. (16)
[…]
$\times\left[\int\limits_{0}^{\infty}\!\!\frac{dx}{e^{\,\beta x}+1}\frac{(x+\omega)^{2}}{F_{3}(x)}-\!\!\int\limits_{\omega+v_{F}q}^{\infty}\!\!\!\!\frac{dx}{e^{\,\beta x}+1}\frac{(x-\omega)^{2}}{F_{4}(x)}\right.\left.\left.+\!\!\int\limits_{0}^{v_{F}q-\omega}\!\!\!\!\frac{dx}{e^{\,\beta x}+1}\frac{(x-\omega)^{2}}{F_{4}(x)}\right]\right\}$ (38)
and
${\rm Im}\,\varepsilon^{\rm T}(\omega,\mbox{$q$})=\frac{\alpha c}{2v_{F}^{2}q\omega^{2}}\sqrt{\omega^{2}-v_{F}^{2}q^{2}}\left[\pi v_{F}^{2}q^{2}-4\!\!\int\limits_{-v_{F}q}^{v_{F}q}\!\!\!\!\frac{dx}{e^{\beta(\omega+x)}+1}\frac{x^{2}}{\sqrt{v_{F}^{2}q^{2}-x^{2}}}\right].$ (39)
We consider first ${\rm Re}\,\varepsilon^{\rm T}$ in the limiting case $\omega\to\infty$. The second integral on the r.h.s. of Eq. (38) vanishes and the remaining two can be rearranged to
$\hskip-11.38109pt-\frac{\sqrt{\omega^{2}-v_{F}^{2}q^{2}}}{\omega^{2}}\!\!\int\limits_{0}^{\infty}\!\!\!\frac{dx}{e^{\,\beta x}+1}\left[\frac{(x+\omega)^{2}}{F_{3}(x)}+\frac{(x-\omega)^{2}}{F_{4}(x)}\right]$
$\hskip-11.38109pt=-\frac{\sqrt{\omega^{2}-v_{F}^{2}q^{2}}}{\omega^{2}}\!\!\int\limits_{0}^{\infty}\!\!\!\frac{dx}{e^{\,\beta x}+1}\left[F_{3}(x)+\frac{v_{F}^{2}q^{2}}{F_{3}(x)}+F_{4}(x)+\frac{v_{F}^{2}q^{2}}{F_{4}(x)}\right]$
$\hskip-11.38109pt\approx-\frac{\omega}{\omega^{2}}\!\!\int\limits_{0}^{\infty}\!\!\!\frac{dx}{e^{\,\beta x}+1}(\omega+\omega)=-2\!\!\int\limits_{0}^{\infty}\!\!\!\frac{dx}{e^{\,\beta x}+1}=-\frac{2\ln 2}{\beta}.$ (40)
Thus, the first term in the figure brackets of Eq. (38) is canceled by Eq. (40) and we obtain
$\lim\limits_{\omega\to\infty}{\rm Re}\,\varepsilon^{\rm T}(\omega,\mbox{$q$})=1,\quad\lim\limits_{\omega\to\infty}{\rm Im}\,\varepsilon^{\rm T}(\omega,\mbox{$q$})=0$ (41)
[the latter equality is evident from Eq.  (39)].
From Eq. (39) it is also seen that ${\rm Im}\,\varepsilon^{\rm T}>0$. This is because the maximum value of the subtracted integral is again reached at $\beta=0$ ($T=\infty$). In this case the integral subtracted in Eq. (39) is
$\max\!\!\int\limits_{-v_{F}q}^{v_{F}q}\!\!\frac{x^{2}dx}{[e^{\beta(\omega+x)}+1]\sqrt{v_{F}^{2}q^{2}-x^{2}}}\!=\frac{1}{2}\!\!\int\limits_{-v_{F}q}^{v_{F}q}\!\!\!\frac{x^{2}dx}{\sqrt{v_{F}^{2}q^{2}-x^{2}}}\!=\frac{\pi}{4}v_{F}^{2}q^{2},$ (42)
which cancels the first term. Thus, at any $T<\infty$, the inequality ${\rm Im}\,\varepsilon^{\rm T}>0$ holds.
In the region $\omega>v_{F}q$ in Figures 3 and 4, $|{\rm Re}\,\varepsilon^{\rm T}|$ and ${\rm Im}\,\varepsilon^{\rm T}$, respectively, given by Eqs. (38) and (39) are shown as the functions of frequency. With increasing $\omega$ from $v_{F}q$ to $1.52\times 10^{12}$ rad/s, ${\rm Re}\,\varepsilon^{\rm T}$ varies to zero remaining negative and then changes its sign and goes to unity. For $\omega>v_{F}q$, ${\rm Im}\,\varepsilon^{\rm T}$ increases from zero and then decreases to zero at $\omega=\infty$. Figures 1 – 4 demonstrate the presence of a threshold at $\omega=v_{F}q$ 67 ; 71 .
At $T=0$, below the threshold, ${\rm Re}\,\varepsilon^{\rm{L,T}}$ are given by the first two terms in Eqs. (18) and (21), whereas ${\rm Im}\,\varepsilon^{\rm{L,T}}=0$. Above the threshold, ${\rm Re}\,\varepsilon^{\rm{L,T}}=1$ and ${\rm Im}\,\varepsilon^{\rm{L,T}}$ are given by the first terms in Eqs. (33) and (39). In each case, an order in the limiting transitions of $\omega$ and $q$ to zero is fixed. At the point of threshold, the derivatives become discontinuous.
The dielectric functions of graphene expressed via the polarization tensor are, by construction, analytic in the upper half-plane of complex frequency and, thus, satisfy the Kramers-Kronig relations. The permittivity $\varepsilon^{{\rm L}}$ is regular at zero frequency, satisfies the standard Kramers-Kronig relations valid for dielectric materials 37 , but has a threshold at $\omega=v_{F}q$. As to the permittivity $\varepsilon^{{\rm T}}$, it is of the most nonconventional character because, according to Eqs. (26) and (30), at nonzero temperature both the real and imaginary parts of $\varepsilon^{{\rm T}}$ have the single pole at $\omega=0$, whereas ${\rm Re}\varepsilon^{{\rm T}}$ also has the double pole. As was noted above, both Eqs.  (26) and (30) were obtained under the condition $T\neq 0$ and it is not possible to consider the limit of zero $T$ in these expressions. At $T=0$ there is no single pole in $\varepsilon^{{\rm T}}$ 71 .
The presence of a single pole, like that one in the imaginary part of the dielectric permittivity of the Drude model, gives rise to the well known additional term in the Kramers-Kronig relations 37 . Similar term appears in the case of a double pole (see Ref. 71 for details). In Ref. 71 it is also shown that the branch points that are present in both $\varepsilon^{{\rm L}}$ and $\varepsilon^{{\rm T}}$ at $\omega=v_{F}q$ do not affect on the form of Kramers-Kronig relations.
As noted in Section 1, all derivations in this article are made for the case of a pristine graphene possessing the zero mass gap parameter $\Delta=2mv_{F}^{2}$, where $m$ is the quasiparticle mass, and chemical potential $\mu$. In the case that $\Delta$ and $\mu$ are not equal to zero, the low-frequency behavior of the dielectric functions of graphene depends of their values preserving the same pole structure as for a pristine graphene. Using expressions for the polarization tensor with the arbitrary values of $\Delta$ and $\mu$ (see, for instance, Ref. 58 ), it is easily seen that if $\Delta$, $\mu$ and $\omega$ simultaneously go to zero one returns back to Eqs. (20), (26) and (30) independently of the order of limiting transitions.
V Discussion
In this article, we investigated the dependence of the dielectric functions of graphene on frequency. The most interesting unusual analytic properties were found for the real part of the transverse function, $\varepsilon^{{\rm T}}$, at low frequencies. Thus, both the real and imaginary parts of $\varepsilon^{{\rm T}}$ possess the single pole at zero frequency. What is more, the spatially nonlocal term in its real part also possesses the double pole, which is not the case for conventional materials according to present views. The double pole should be also present in the response functions of other 2D Dirac materials such as germanene 71.1 ; 71.2 ; 71.3 , silicene 71.4 ; 71.5 ; 71.6 , phosphorene 71.7 ; 71.8 ; 71.9 , and stanene 71.10 ; 71.11 ; 71.12 . In spite of the presence of a double pole, the dielectric functions of graphene satisfy all the physical demands considered above. They possess the positive imaginary parts, which describe dissipation on the basis of first principles, and satisfy the Kramers-Kronig relations expressing the condition of causality. Because of this, an attempt 66 to modify the polarization tensor in order to remove the double pole predicted by the first principles of quantum field theory is unjustified.
There is also a long-standing problem called the Casimir puzzle. To bring the theoretical predictions of the fundamental Lifshitz theory in agreement with the measurement data, the dielectric response of metals at low frequencies was described by the plasma model possessing the double pole at $\omega=0$ (see Refs. 72 ; 73 ; 74 ; 75 for a review). However, as mentioned above, this model is applicable only at high frequencies. That is why an example of graphene, whose dielectric function possessing the double pole at zero frequency is derived starting from first physical principles and leads to an agreement with measurements of the Casimir force, may pave the way for resolution of the Casimir puzzle.
VI Conclusions
In the foregoing, we listed several phenomenological theoretical approaches used for investigation of the dielectric response of graphene. It is underlined that at the characteristic energies below approximately 3 eV the spatially nonlocal response functions of graphene can be derived within the Dirac model starting from first principles of thermal quantum field theory. The obtained dielectric functions are useful for a theoretical description of many physical phenomena in graphene systems, such as the Casimir and Casimir-Polder forces both in equilibrium situations and out of thermal equilibrium, radiative heat transfer, atomic friction, surface plasmons etc.
According to our results, these functions possess all the properties necessary for the dielectric functions and their transverse part has the double pole at zero frequency at any nonzero wave vector. The above discussion allows to make a conjecture that the spatially nonlocal transverse electric response function of metals possesses the double pole in the region of evanescent waves like it holds for graphene. Recently it was demonstrated 76 that the predictions of classical electrodynamics using the Drude dielectric function for the field of oscillating magnetic dipole reflected from a copper plate, which is fully determined by the transverse electric evanescent waves, are in contradiction with the measurement data. Future progress in investigation of such physical phenomena as the Casimir effect, atomic friction, radiative heat transfer, near-field optical microscopy, total internal reflection and frustrated total internal reflection is closely allied to the resolution of this problem.
This work was supported by the State Assignment for Basic Research (project FSEG-2026-0018).
Appendix A The list of auxiliary functions used in the article
Here, for the readers convenience, we present a summary of the auxiliary functions and other notations used in this article (see Table A1).
Table 1: The definitions of auxiliary functions and other notations.
$\alpha$ $\frac{e^{2}}{\hbar c}$ is the fine structure constant
$\beta$ $\frac{\hbar}{2k_{B}T}$
$\gamma$ $v_{F}q\beta$
$\kappa$ $\frac{\omega}{v_{F}q}$
$B_{1}$ $2\int\limits_{0}^{1}\!\!\!\frac{t^{2}dt}{\sqrt{1-t^{2}}}\frac{e^{\gamma t}}{(e^{\gamma t}+1)^{2}}$
$B_{2}$ $2\int\limits_{1}^{\infty}\!\!\!\frac{t^{2}dt}{\sqrt{t^{2}-1}}\frac{e^{\gamma t}}{(e^{\gamma t}+1)^{2}}$
$F_{1,2}(x)$ $\sqrt{v_{F}^{2}q^{2}-(x\pm\omega)^{2}}$
$F_{3,4}(x)$ $\sqrt{(x\pm\omega)^{2}-v_{F}^{2}q^{2}}$
$I_{1}$ $\int\limits_{\kappa}^{1}\!\!\frac{t^{2}dt}{\sqrt{1-t^{2}}}\left(\frac{1}{e^{\gamma t}e^{-\gamma\kappa}+1}-\frac{1}{e^{\gamma t}e^{\gamma\kappa}+1}\right)$
$I_{2}$ $\int\limits_{-\kappa}^{\kappa}\frac{dt}{e^{\gamma t}e^{\gamma\kappa}+1}\,\frac{t^{2}}{\sqrt{1-t^{2}}}$